# From Efficiency Gains to Rebound Effects: The Problem of Jevons’ Paradox in AI’s Polarized Environmental Debate

### v1

@@ 3.2. Economic Rebound Effects
Direct economic rebound effects refer to situations in which the improved efficiency of a product decreases its price, therefore leading to an increase in its consumption. In the centuries since the concept of Jevons’ paradox was developed, similar rebound effects have been observed with respect to energy (Giampietro and Mayumi, 2018) and water (Dumont et al., 2013), as well as in domains such as road travel (where improvements to roads have been found to result in increased congestion (Thomson, 1972)) and agriculture (where increasing the yield of a crop makes it more profitable to grow it, thereby increasing land use overall (Barbieri et al., 2019)). They are also the most common type of indirect effects that are discussed in conjunction with many technological innovations (see (Kaack et al., 2022; Willenbacher et al., 2021)). AI is no exception to direct economic rebound effects. While efficiency improvements are being made to the hardware used for training and deploying AI models (NVIDIA, 2024; Patterson et al., 2022a; Baidu, 2023), NVIDIA shipped 3.7 million GPUs in 2024 (more than a million more units than in 2023) due to increased demand, despite these improvements in efficiency (Shah, 2024). The data centers that host this hardware are also becoming more efficient, with the average PUE (Power Usage Efficiency) dropping steadily in recent years (Evans and Gao, 2016; Sharma et al., 2015), the energy use and environmental impacts of these data centers has been rising (as described in Section 2).
@@
On a more macro level still, economy-wide rebound effects occur when an innovation provokes far-reaching changes in the production and use of other goods, producing flow-on effects in the economy at large. For instance, improvements in global energy efficiency and fuel use have enabled the development of new economies on a global scale, allowing the creation of new industries (Özsoy, 2024). These types of rebound effects can be particularly large for so-called “general-purpose technologies” such as electricity, the steam engine and the Internet, since these can have global spillover effects on industries and labor markets (Sorrell, 2009; Sorrell et al., 2007). Estimating the scope of economy-wide rebound effects has historically been difficult (Stern, 2020), especially for technologies (Gossart, 2015). One way in which AI is concretely impacting the economy at large is via PPAs, which represent long-term contracts between electricity producers and customers that promise to purchase the energy that they generate over a horizon that can span up to 20 years, which enables the growth and development of renewable energy. In 2020 alone, the biggest tech companies (Amazon, Microsoft, Meta, Google) accounted for almost 30% of all PPAs purchased by corporations worldwide (Varro and Kamiya, 5). This can also have ripple effects on the global economy as a whole, lowering the risk of building out new renewable energy generation infrastructure, which typically requires a large up front investment that can be hard for smaller operators to contribute (Tantau and Niculescu, 2022). Also, while part of the resulting power generated by the producer goes to the companies, the remainder is added to the energy grid in that region, allowing other users to profit from this additional energy to build out new buildings and businesses. It has also been proposed that foundation models are, in fact general purpose technologies with the potential to impact the whole economy, as well as labor markets (Eloundou et al., 2023) — analyzing the impacts of these changes in the future will be necessary to validate this claim (we will discuss approaches to tracking AI’s indirect impacts in Section 3.4).
@@ 3.3. Societal and Behavioral Rebound Effects
As users of technology and citizens of society, our behaviors are shaped and affected by the technologies that we use. This has increasingly become the case as technologies are integrated into our daily lives and activities such as shopping, travel, and household chores. In the current section, we explore the ways in which AI can impact our behaviors and how these changes, in turn, can have broader environmental impacts.
@@
Induction effects occur when the savings in resource consumption (e.g. energy or water) gained through improved efficiency are exceeded by the increased consumption and thereby production of either the materials or the final products. For instance, the advent of fast fashion and mass-produced furniture have reduced the quantity of natural resources – such as wood or cotton – needed to produce a single piece of clothing or furniture. However, the amount of plastics, petrochemicals and waste that are engendered by the increased accessibility of these items and therefore their increased consumption are damaging to the environment (Niinimäki et al., 2020). In the case of AI, a prime example of this is targeted advertising, one of the most lucrative commercial applications of AI with revenues reaching \$36 billion in 2024 (Statista, 2024). It also constitutes the main source of revenue for many technology companies such as Google and Meta (Otto, Melissa, 2024); Amazon’s AI-powered product recommendation tool generates almost a third of its annual sales (MacKenzie, 2013). While it is hard to obtain quantitative information on the impact of AI on e-commerce given the corporate secrecy that surrounds it, recommendation and ranking is an active topic of research in the AI community given the quantity of user data that exists and continues to be gathered in our interactions with digital platforms (Chen et al., 2009; Dominowska and Josifovski, 2008; Choi and Lim, 2020). In fact, the rapid expansion of online advertising and the direct and indirect environmental impacts of this expansion have been emphasized in research papers and scientific reports over the last years (Hartmann et al., 2023; PlanetTracker, 2024). The contribution of AI to this expansion is worth considering when measuring the performance of new algorithms and approaches (Dauvergne, 2022), especially given the net-zero and emission reduction goals of many AI-driven companies whose revenue models depend on it (Willenbacher et al., 2021).
@@
Time rebound effects occur when an innovation changes consumers’ use of their time, which then frees up (or removes) time for other activities that they carry out. For example, if a vacuum cleaner reduces the time needed to clean the floors in a house, it can free up time for leisure activities such as reading or sports. The sustainability benefits of these time savings are far from straightforward, because what we do with this newly-freed time can have negative environmental impacts as well – if time otherwise spent doing chores is now spent shopping or traveling, this can increase our overall carbon footprint (Buhl and Acosta, 2016). AI is undeniably impacting how we spend our time both at work and at home: there are multiple activities that have already been streamlined using AI in the last decade – such as AI-enabled navigation that lets us avoid traffic jams, or toll booths that automatically read license plates and send us invoices in the mail, and robots that save us time by sweeping our floor. Initial work on the environmental consequences of this shift has proposed that AI-enabled automation in household tasks and in the workplace actually leads to more resource consumption and higher environmental impacts overall (Ertel and Bonenberger, 2024). This contradicts research carried out on AI navigation specifically, which has been shown to reduce emissions by an average of 3.4%, based on actual customer data gathered by Google (Arora et al., 2021). This highlights the difficulty of making far-reaching claims about the sustainability impacts of a given innovation without considering the spillover effects and behavioral changes that this may engender more broadly– we address these challenges in the following section.
@@ 3.4. Tracking and Mitigating AI’s Rebound Effects
The main challenges with measuring and mitigating indirect impacts and rebound effects is their uncertainty and heterogeneity. While tracking direct impacts can be done via a single metric, be it liters of water or tons of carbon dioxide emitted, rebound effects can engender social, economic, and behavioral impacts in different areas of society (Sovacool et al., 2015). This is why it helps to look at a multitude of complementary approaches that can help the AI community go beyond the binary narrative towards more holistic, and effective, ways of analyzing and addressing these impacts. We discuss some of these approaches below.
@@
Starting with the material rebound effects of digital technologies such as AI, it can be tempting to make comparisons such as the research described in Section 3.1, which carry out LCA analysis comparisons between, e.g a book and an e-reader. However, we are currently lacking tools and approaches that would allow us to meaningfully compare the environmental impacts of humans versus the same tasks performed by humans alone or assisted by AI. Initial work on this subject has concluded that “AI systems emit between 130 and 1500 times less CO2e per page of text generated compared to human writers, while AI illustration systems emit between 310 and 2900 times less CO2e per image than their human counterparts,” however, the authors themselves note that their analyses “do not account for social impacts such as professional displacement, legality, and rebound effects” (Tomlinson et al., 2024). More qualitative and multi-faceted analyses, such as those that take into account aspects such as the length of time that content will actually spend existing in the world can help shed light on the issue, as well as more empirical research to study how human artists and writers use AI technologies and how that impacts the environment.
@@
One of the difficulties in tracking AI’s economic rebound effects stems from the integration of AI into many existing products and services, making it difficult to disentangle growth inherent in the field with the growth influenced by the development of AI. Gathering more information about the impacts of AI tools and developing frameworks to measure their impacts will become increasingly important as AI becomes a mainstream technology. In the case of induction effects in particular, establishing a baseline across industries – for instance of the environmental impacts of technologies that are currently in use – and using it to measure the impact of AI can be a good way of tracking longitudinal trends. Requiring companies using AI in applications such as targeted advertising to indicate this usage, as well as what data was used to choose an ad for a specific product, can help users better understand why they are receiving ads and contribute towards a more mindful consumption (Dangelico and Vocalelli, 2017; Minton et al., 2012).
@@
While focusing on efficiency alone has largely been sufficient in the past to curb the environmental impacts of rising compute (Masanet et al., 2020b), given the astronomical rise in compute needed for AI services since the generative AI revolution, complementary efforts are needed to address these impacts (Shehabi et al., 2024). Historically, economic policies that increase cost of usage of the initial resources have been shown to help control these kinds of rebound effects (Freire-González and Puig-Ventosa, 2015). Initiatives such as energy efficiency certifications (e.g. Energy Star, Energy Score) can also reduce resource use by incentivizing the usage of more efficient tools and systems for a set of standardized tools, such as appliances (Datta and Filippini, 2016). A related proposal has been introduced for AI models (Luccioni et al., 2024), although the lack of centralized authority for regulating AI (either at a national or international level) makes it difficult to operationalize such a proposal. Establishing regulation that would require companies to disclose the usage of AI in environmentally harmful applications such as oil and gas exploration, which is currently not reflected in company documents.
@@
Increasingly compact and efficient computing devices that require less natural resources than previous generations can be seen as an obvious gain in terms of environmental impacts, and many different hardware types continue to be developed as an alternative to traditional GPUs and CPUs (Jouppi et al., 2017; Jia et al., 2019). Also, new approaches to distributed and decentralized AI model training can help redistribute the load and make it possible to train and finetune models across datacenters and continents, either by carrying out full decentralized training (e.g.  (Jaghouar et al., 2024)) or by redirecting training processes depending on which cloud region has the lowest carbon intensity at a given time of day (Dodge et al., 2022). Above and beyond sustainability considerations, adopting more adaptable hardware and more flexible training approaches can help lower the barrier to entry into the field of AI and push back against industry monopolization of compute and power (Abdalla and Abdalla, 2021; Abdalla et al., 2023).

### v2

@@ 3.2. Economic Rebound Effects
Direct economic rebound effects refer to situations in which the improved efficiency of a product decreases its price, therefore leading to an increase in its consumption. In the centuries since the concept of Jevons’ paradox was developed, similar rebound effects have been observed with respect to energy (Giampietro and Mayumi, 2018) and water (Dumont et al., 2013), as well as in domains such as road travel (where improvements to roads have been found to result in increased congestion (Thomson, 1972)) and agriculture (where increasing the yield of a crop makes it more profitable to grow it, thereby increasing land use overall (Barbieri et al., 2019)). They are also the most common type of indirect effects that are discussed in conjunction with many technological innovations (see (Kaack et al., 2022; Willenbacher et al., 2021)). AI is no exception to direct economic rebound effects. While efficiency improvements are being made to the hardware used for training and deploying AI models (NVIDIA, 2024; Patterson et al., 2022a; Baidu, 2023), NVIDIA shipped 3.7 million GPUs in 2024 (more than a million more units than in 2023) due to increased demand, despite these improvements in efficiency (Shah, 2024). The data centers that host this hardware are also becoming more efficient, with the average PUE (Power Usage Efficiency) dropping steadily in recent years (Evans and Gao, 2016; Sharma et al., 2015); however, the energy use and environmental impacts of these data centers has been rising (as described in Section 2).
@@
On a more macro level still, economy-wide rebound effects occur when an innovation provokes far-reaching changes in the production and use of other goods, producing flow-on effects in the economy at large. For instance, improvements in global energy efficiency and fuel use have enabled the development of new economies on a global scale, allowing the creation of new industries (Özsoy, 2024). These types of rebound effects can be particularly large for so-called “general-purpose technologies” such as electricity, the steam engine and the Internet, since these can have global spillover effects on industries and labor markets (Sorrell, 2009; Sorrell et al., 2007). Estimating the scope of economy-wide rebound effects has historically been difficult (Stern, 2020), especially for technologies (Gossart, 2015). One way in which AI is concretely impacting the economy at large is via PPAs, as mentioned in Section 2.2. In 2020, Amazon, Microsoft, Meta, and Google alone accounted for almost 30% of all PPAs purchased by corporations worldwide (Varro and Kamiya, 5), changing the scope and extent of the mechanism as a whole. Renewable energy PPAs mitigate the risk of investment in renewable energy infrastructure, thereby supporting its development (Tantau and Niculescu, 2022). Further, while a portion of the generated renewable energy is transmitted to purchasers in practice, the remainder is added to the energy grid in that region, allowing other users to benefit from its availability. It has also been proposed that foundation models are general purpose technologies with corresponding potential for broad impacts to the economy and labor markets (Eloundou et al., 2023). Analyzing the impact of these changes in the future will be necessary to validate this claim (we discuss approaches to tracking AI’s indirect impacts in Section 3.4).
@@ 3.3. Societal and Behavioral Rebound Effects
As users of technology and citizens of society, our behaviors are shaped and affected by the technologies we use. This has increasingly become the case for AI as it becomes more intertwined in daily activities such as shopping, travel, household chores, and the workplace. In this section, we describe ways in which AI can impact human behaviors and how those changes, can have broader environmental impacts in turn.
@@
Induction effects occur when the savings in resource consumption (e.g. energy or water) gained through improved efficiency are exceeded by the increased consumption and thereby production of either the materials or the final products. For instance, the advent of fast fashion and mass-produced furniture have reduced the quantity of natural resources, such as wood or cotton, needed to produce a single piece of clothing or furniture. However, the amount of plastics, petrochemicals and waste that are engendered by the increased accessibility of these items and therefore their increased consumption are damaging to the environment (Niinimäki et al., 2020). In the case of AI, a prime example of this is targeted advertising, one of the most lucrative commercial applications of AI, with revenues reaching \$36 billion in 2024 (Statista, 2024). Advertising also constitutes the main source of revenue for many technology companies such as Google and Meta (Otto, Melissa, 2024); Amazon’s AI-powered product recommendation tool generates almost a third of its annual sales (MacKenzie, 2013). Although it is difficult to obtain quantitative data on the impact of AI on e-commerce, the details of which are largely considered sensitive trade secrets, the corresponding machine learning approaches of recommendation and ranking are active topics of research in the AI community, powered by an enormous quantity of user data being continuously gathered through interactions with digital platforms (Chen et al., 2009; Dominowska and Josifovski, 2008; Choi and Lim, 2020). The rapid expansion of online advertising its direct and indirect environmental impacts have been emphasized in research papers and scientific reports over the last years (Hartmann et al., 2023; PlanetTracker, 2024). The contribution of AI to this expansion is worth considering when measuring the performance of new algorithms and approaches (Dauvergne, 2022), especially given the net-zero and emission reduction goals of many AI-driven companies whose revenue models depend on it (Willenbacher et al., 2021).
@@
Time rebound effects occur when an innovation changes consumers’ use of their time, which then frees up (or removes) time for other activities that they carry out. For example, if a vacuum cleaner reduces the time needed to clean the floors in a house, it can free up time for leisure activities such as reading or sports. Measuring the sustainability benefits of these time savings is far from straightforward, because what we do with this newly-freed time can have negative environmental impacts as well, e.g. if time otherwise spent doing chores is now spent shopping or traveling, this can increase one’s overall carbon footprint (Buhl and Acosta, 2016). AI is undeniably impacting how time is spent at work and at home: many activities have already been streamlined using AI over the last decade, such as AI-enabled navigation that avoids traffic congestion, automated license plate recognition and tolling on highways, and robotic vacuums. Initial work on the environmental consequences of these shifts has proposed that AI-enabled automation in household tasks and in the workplace actually leads to greater resource consumption and higher negative environmental impacts overall (Ertel and Bonenberger, 2024). This contradicts research carried out on AI-assisted navigation specifically, which has been shown to reduce emissions by an average of 3.4%, based on data gathered by Google on users of Google Maps (Arora et al., 2021). This discrepancy highlights the difficulty of making far-reaching claims regarding the sustainability impacts of a given innovation without considering the spillover effects and behavioral changes that it may engender more broadly.
Indirect policy effects. In addition to societal and behavioral forces, there are also indirect geopolitical and policy effects. International policy interventions, particularly those aiming to bolster national self-sufficiency from global supply chains, can provoke unintended increases in environmental burdens through mineral demand amplification, industrial deregulation, and competitive overproduction. The dynamics of Koomey’s Law, by which the energy efficiency of computing has historically doubled approximately every 1.57 years, can also be affected by geopolitical disputes and trade wars. If chip sales are restricted, for example, nations have to make do with less powerful and less energy-efficient semiconductors. As national strategies to monopolize or re-shore semiconductor capacity develop, this can further accelerate national rivalries and result in the replication of systems and further resource-intensive scaling of model training infrastructure. The United States CHIPS and Science Act (2022) is one example. This legislation aims to re-shore semiconductor fabrication. While its primary goal is to reduce dependency on Chinese manufacturing, it has the potential to intensify domestic resource demands. Semiconductor fabrication facilities (known as fabs) impose massive environmental burdens, consuming large quantities of energy, water and specialized chemicals while generating substantial carbon emissions (Villard, 2015). The domestic expansion of fabs in the United States shifts rather than reduces these environmental impacts, potentially introducing new ecological pressures in regions with water scarcity issues like Arizona, where several new facilities were planned. Meanwhile the ongoing resource pressures continue in China and Taiwan. Moreover, the race to decouple US semiconductor supply chains from geopolitical adversaries has led to intensified extraction of rare-earth elements, lithium, cobalt, and high-purity quartz. This geopolitical reorganization of material flows extends the environmental footprint of AI beyond traditional carbon accounting. China restricting the export of several rare earth elements has meant that mines in Ukraine and Australia are already under expanded pressure to meet US demand. The increased competition for critical minerals has both social and ecological impacts, and risks even greater environmental harms in conflict-affected regions.
These indirect policy effects risks are further compounded when environmental deregulation coincides with large-scale investment in AI-related infrastructure. The Trump administration in 2025 has been marked by a comprehensive rollback of environmental protections, including a withdrawal from the Paris agreement, significantly altering the United States’ approach to climate and energy policy. Simultaneously, President Trump announced a \$500 billion USD investment in AI infrastructure known as Project Stargate, which aims to significantly increase the number of very large (hyperscale) data centers and bypass environmental impact assessments while offering public subsidies. These policy shifts can lock-in energy and environmental demands that outpace any computational efficiency gains.
This decoupling of AI infrastructure from environmental responsibility illustrates a critical rebound mechanism: technical efficiency improvements, such as high-efficiency datacenter designs or AI model efficiency optimizations, are insufficient when broader regulatory and policy conditions incentivize unchecked growth. By facilitating rapid deployment without ecological constraint, these conditions create an emissions burden with resource demands that scale with model proliferation. In such contexts, even the best-case efficiency improvements under Koomey’s Law may be dwarfed by the sheer increase in volume of compute infrastructure. Then the indirect effects of these policies on the environment are no longer marginal, they become structural. The recent DeepSeek model is actually an interesting addition to the debate around efficiency and energy because, on the one hand, it was trained using relatively less energy and compute compared to previous generations of language models (due mostly to the export limitations on GPUs to China, which limited the amount of chips that was available for training it), but on the other hand, this was only possible by leveraging synthetic data generated by these models. Also, the model itself is very large –almost 700B parameters for the R1 model – meaning that deploying it requires access to multiple GPUs with several hundreds of GB of memory, which is significantly more than smaller models of a smaller size, which can contribute to more energy use overall as organizations deploy the model in user-facing applications. Finally, as detailed in the DeepSeek-R1-Zero report (DeepSeek-AI et al., 2025), the model requires much more inference-time computation and energy than previous approaches due to its reasoning abilities. This results in the generation of much longer answers to queries detailing the steps that the model is going through to provide the final response, which leads to more inference-time compute and energy demands.
@@ 3.4. Tracking and Mitigating AI’s Rebound Effects
The main challenges in measuring and mitigating indirect impacts and rebound effects are their uncertainty and heterogeneity. While tracking direct impacts can largely be achieved by monitoring a finite set of relatively well-defined metrics, such as liters of water consumed or tons of CO2 emitted, rebound effects by definition encompass social, economic, and behavioral impacts across different areas of society (Sovacool et al., 2015). Realizing effective progress in characterizing AI’s rebound effects necessitates moving beyond the binary narrative towards more holistic assessment that incorporates a variety of complementary approaches. We discuss some of these approaches below.
@@
Starting with the material rebound effects of digital technologies such as AI, it can be tempting to make comparisons such as the research described in Section 3.1, which carry out LCA comparisons between e.g. a book and an e-reader. However, we are currently lacking sufficient methodology and data that would allow for meaningful comparison of the environmental impacts of humans versus the same tasks performed by humans alone or assisted by AI. Initial work on this subject has concluded that “AI systems emit between 130 and 1500 times less CO2e per page of text generated compared to human writers, while AI illustration systems emit between 310 and 2900 times less CO2e per image than their human counterparts,” however, the authors themselves note that their analyses “do not account for social impacts such as professional displacement, legality, and rebound effects” (Tomlinson et al., 2024). More qualitative and multi-faceted analyses are needed, such as those that take into account aspects like the length of time that content will actually be useful to readers and the way in which the data was collected, can help shed light on the issue, as well as more empirical research to study how human artists and writers use AI technologies and how that impacts the environment.
@@
A key challenge to tracking AI’s economic rebound effects stems from the severe scarcity of information regarding the nature and extent of AI’s actual permeation throughout the economy. The rapid integration of AI into many existing products and services, combined with vague and often broad definitions of AI, has made it difficult to disentangle AI-influenced growth from growth that might have occurred regardless. This challenge is intensified by the perception that details on corporate AI model use represent valuable intellectual property and trade secrets. Gathering more information about specific AI deployment scenarios, i.e. which AI models are actually being deployed, in which sectors, what are they being used for in practice, and what are the impacts of those uses, will become increasingly important as AI becomes established as a mainstream technology. Connecting data on AI model resource consumption to corresponding positive or negative impacts arising from those models is key to understanding tradeoffs and informing decision-making. In the case of induction effects in particular, establishing baselines across industries, for instance documenting the environmental impacts of technologies currently in use, and measuring changes to those baselines as AI is integrated would enable tracking longitudinal trends. Requiring companies using AI in applications such as targeted advertising to indicate this usage, as well as what data was used to choose an ad for a specific product, can help users better understand why they are receiving ads and contribute towards more mindful consumption (Dangelico and Vocalelli, 2017; Minton et al., 2012).
@@
While focusing on efficiency alone has largely been sufficient in the past to curb the environmental impacts of rising compute (Masanet et al., 2020b), given the astronomical rise in computational (and corresponding energy and natural) resources needed to power AI services resulting from generative AI, complementary efforts are needed to address these impacts (Shehabi et al., 2024). Historically, economic policies that increase the cost of these resources have been shown to help control certain rebound effects (Freire-González and Puig-Ventosa, 2015). Initiatives such as energy efficiency certifications (e.g. the U.S EPA’s Energy Star certification, Energy Labels in the EU) can also reduce resource use by incentivizing the sales and production of more efficient tools and systems, successful for increasing energy efficiency for household appliances (Datta and Filippini, 2016). Related proposals have been introduced for AI models (Luccioni et al., 2024; Fischer et al., 2022), although the lack of centralized authority for regulating AI (either at a national or international level) has made it difficult so far to operationalize such a proposal. Regulations could also serve an important role in mandating reporting of the necessary data described above, e.g. requiring disclosure of the usage of AI for environmentally harmful applications such as oil and gas exploration, which although highly relevant, is not typically reflected in corporate Environment and Sustainability Governance (ESG) documents.
@@
Resource-efficient computing devices that require less natural resources (energy, water, raw materials) to build and use can be seen as an obvious gain in terms of environmental impacts, and different hardware platforms are being developed towards this end as an alternative to GPUs and CPUs, which represent the current status quo of AI computing hardware (Jouppi et al., 2017; Jia et al., 2019). It is worth noting that there is typically a trade-off between resource efficiency and generality in terms of the computational capabilities of the hardware: substantial efficiency gains typically require the hardware to be correspondingly specialized to certain types of computation, such as a specific family of AI models, which can accelerate indirect effects via lock-in, increasing dependence on a narrow set of vendors or products. New approaches to distributed and decentralized computation, in which AI model workloads can be distributed across machines over wide distances rather than running within a single computing cluster or datacenter, can also help distribute environmental burden geographically in terms of resource demand and local impacts. Examples include fully decentralized training (Jaghouar et al., 2024) or redirecting training processes depending on which cloud region has the lowest carbon intensity at a given time of day (Dodge et al., 2022). Above and beyond sustainability considerations, adopting more flexible AI hardware and algorithms can help lower the barrier to entry into the field of AI and push back against industry monopolization of computational resources, and corresponding power imbalance that this creates (Abdalla and Abdalla, 2021; Abdalla et al., 2023).